OpenStudy (anonymous):

a few questions on circumference if i can get help. will give medal and fan anybody who helps

OpenStudy (johnweldon1993):

What do you need to know?

OpenStudy (anonymous):

what do you have a question on?

OpenStudy (anonymous):

What is the approximate circumference of a semicircle with a radius of 35 inches? A.1923.25 inches B.219.8 inches C.109.9 inches D.54.95 inches

OpenStudy (johnweldon1993):

You know semicircle is half a circle... So just calculate the circumference of a full circle with that radius...and divide it by 2 \[\large C_{\text{semicircle}} = \frac{2 \pi r}{2}\] \[\large C_{\text{semicircle}} = \pi r\]

OpenStudy (anonymous):

i dont know what the whole is of a circle

OpenStudy (anonymous):

i dont know anything about circumference